okay i dont know how to do the whole typy in htingy but it says
"the square root of x+2 = the square root of 3x-6 so is it 4?

Answer by stanbon(75887)

okay i dont know how to do the whole typy in htingy but it says
"the square root of x+2 = the square root of 3x-6 so is it 4?
sqrt(x+2) = sqrt(3x-6)
Square both sides of the equation to get:
x+2 = 3x-6
2x=8
x=4
Check the answer:
sqrt(4+2) = sqrt(3(4)-6)
sqrt(6)= sqrt(6)
It checks.
Cheers,
stan H.